OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 4x^3 - 5x^2 + 2x + 11; k= 3

OpenStudy (anonymous):

My choices were 68, 136, 80, & 24.

OpenStudy (raden):

The remainder of f(x) when divided by (x - k) is f(k) here k = 3, so just find the value of f(3). replace x = 3 into f(x), it means f(3) = 4(3)^3 - 5(3)^2 + 2(3) + 11 = ..... ?

OpenStudy (anonymous):

Oh, really? You can do that? Wait.. When you add that up, it comes to -151... That's not one of my choices.. :(

OpenStudy (isaiah.feynman):

What @RadEn did is called the remainder theorem.

OpenStudy (raden):

f(3) = 4(3)^3 - 5(3)^2 + 2(3) + 11 f(3) = 4(27) - 5(9) + 6 + 11 f(3) = 108 - 45 + 6 + 11 = ... ? now, smack down this! :)

OpenStudy (isaiah.feynman):

Remember BEDMAS!!!

OpenStudy (anonymous):

Hey, I found it. :D ANyways, that's what I get for typing that into my calculator... Lazy Americans. -.- What is BEDMAS?

OpenStudy (anonymous):

?

OpenStudy (isaiah.feynman):

brackets Exponents Division Multiplication Addition Subtraction.